# Shadows of Rotating Black Hole Solution with a Dark Matter Halo

### v1

@@ IV Constraints from EHT observational data
Figure 5: Angular diameter observable $\Theta_{d}$ for the black hole shadows as a function of parameters $r_{s}$ and $\rho_{s}$. The black curves correspond to $\Theta_{d}=41.7\mu as$ (for top panels) and $\Theta_{d}=39\mu as$ (for bottom panels) within $1\sigma$ region of the measured angular diameter, $\Theta_{d}=48.7\pm 7\mu as$, of the SgrASgrA* black hole (top panels) and $\Theta_{d}=42\pm 3\mu as$ of the M87M87* black hole reported by the EHT. The blue dashed curve corresponds to $48.7\mu as$ and $42\mu as$.
@@
The observational angular diameters of the shadow images of the SMBHs M87* and Sgr A* are $\Theta_{d}=42\pm 3\mu as$ and $\Theta_{d}=48.7\pm 7\mu as$, respectively (2019ApJ...875L...1E,Akiyama and et al. (2022) (Event Horizon Telescope Collaboration)). We would like to point out that, for simplicity, we do not take into account the possible uncertainties in the mass and distance measurements of the black holes in the numerical estimations.
@@
Then in the top panel of Fig. 5, we present the density plots of the characteristic scale $r_{s}$ and characteristic density $\rho_{s}$ of the black hole that are fitted with the measured values of the angular diameter of SgrA*, where we assume the mass of SMBH SgrA* as $M\approx 4\times 10^{6}M_{\odot}$ and the distance from Earth to SMBH SgrA* as $d\approx 8kpc$ (Akiyama and et al. (2022) (Event Horizon Telescope Collaboration)), for the different fixed values of the BH spin $a$. To be more informative, we show the lower bound $\Theta_{d}=41.7\mu as$ (black curves on top panels of Fig. 5) and the mean value $\Theta_{d}=48.7\mu as$ (blue dashed lines on top panels of Fig. 5 of the angular diameter of the supermassive black hole SgrA*. It is clear from Fig. 5 that the values of the BH parameters $r_{s}$ and $\rho_{s}$ should be relatively high to coincide with the EHT observational data for SgrA *.
@@
In addition, we performed the same analysis for the measured value of the angular diameter of SMBH M87* (2019ApJ...875L...1E) in the bottom panel of Fig. 5. Again, we present the lower bound $\Theta_{d}=39\mu as$ (solid black lines in the bottom graphics of Fig. 5), and mean value $\Theta_{d}=42$ (blue dashed lines in the bottom panels of Fig. 5) of the angular diameter of the supermassive black hole M87*. Similarly, the characteristic scale $r_{s}$ and the characteristic density $\rho_{s}$ of the rotating BH in the dark matter halo have to be relatively high to coincide with the EHT observational data for M87*.
V Conclusion
@@
The Event Horizon Telescope (EHT) observations of the shadows of supermassive black holes, like Sgr A* at the Milky Way’s center and M87* in the nearby Messier 87 galaxy, provide an exceptional and natural laboratory for studying the gravitational properties of spacetime geometry around black holes, allowing tests in both strong and weak gravitational field regimes. In this work, we have generated a rotating BH space-time in a dark matter halo applying the NJA method to the static spacetime metric and the components of the obtained metric tensor are provided in (8). Then dependence of the horizon radius on the BH spin $a$ is shown in Fig. 2 for different values of the characteristic density $\rho_{s}$ and characteristic scale factor $r_{s}$. It is shown that increasing spacetime parameters $r_{s}$ and $\rho_{s}$ causes an enlarging the horizon radius of the rotating BH in a dark matter halo.
@@
We have also studied the shadow cast by the spinning axially symmetric black hole surrounded by a Dehnen-type dark matter halo. Using the Hamilton-Jacobi equation, we found photon motion in the vicinity of the rotating BH in dark matter halo, which enabled us to find photon sphere $r_{ph}$ and the shadow radius. The results obtained showed that the presence of the dark-matter halo causes an enlarged shadow presented in Fig. 3.
@@
Additionally, we have analyzed the photon sphere and shadow radius of the rotating BH in dark matter halo by giving numerical values of the photon sphere and shadow radius in Fig. 4.
@@
To strengthen our pure theoretical analysis, we got constraints for BH parameters $\rho_{s}$ and $r_{s}$ from EHT observational data for spacetime parameters of supermassive black holes SgrA* and M87* in Sec. IV. Our provided analysis for the angular diameter $\Theta_{d}$ demonstrates that the characteristic density $\rho_{s}$ and the characteristic scale factor $r_{s}$ of the rotating BH in the Dehnen-type dark matter halo should be relatively high to fit the EHT observed data of the SMBHs images (look to Fig.5).
One may conclude from the performed study that the influence of a Dehnen-type dark matter halo in the environment of supermassive black holes SgrA* and M87* can not be excluded. More precise data from the EHT collaboration is required in order to estimate possible presence of dark matter in the centers of the Milky Way and Messier 87 galaxies.
Acknowledgements

### v2

@@ 5 Constraints from EHT observational data
Figure 7: Angular diameter observable $\Theta_{d}$ for the black hole shadows as a function of parameters $r_{s}$ and $\rho_{s}$. The black curves correspond to $\Theta_{d}=41.7\mu as$ (for top panels) and $\Theta_{d}=39\mu as$ (for bottom panels) within $1\sigma$ region of the measured angular diameter, $\Theta_{d}=48.7\pm 7\mu as$, of the SgrASgrA* black hole (top panels) and $\Theta_{d}=42\pm 3\mu as$ of the M87M87* black hole reported by the EHT. The blue dashed curve corresponds to $48.7\mu as$ and $42\mu as$, respectively.
@@
The observational angular diameters of the shadow images of the SMBHs M87* and Sgr A* are $\Theta_{d}=42\pm 3\mu as$ and $\Theta_{d}=48.7\pm 7\mu as$, respectively (Event Horizon Telescope Collaboration et al. [2019],Akiyama and et al. (2022) [Event Horizon Telescope Collaboration]). We would like to point out that, for simplicity, we do not take into account the possible uncertainties in the mass and distance measurements of the black holes in the numerical estimations.
@@
Then in the top panel of Fig. 7, we present the density plots of the characteristic scale $r_{s}$ and characteristic density $\rho_{s}$ of the black hole that are fitted with the measured values of the angular diameter of Sgr A*, where we assume the mass of SMBH SgrA* as $M\approx 4\times 10^{6}M_{\odot}$ and the distance from Earth to SMBH Sgr A* as $d\approx 8kpc$ (Akiyama and et al. (2022) [Event Horizon Telescope Collaboration]), for the different fixed values of the BH spin $a$. To be more informative, we show the lower bound $\Theta_{d}=41.7\mu as$ (black curves on top panels of Fig. 7) and the mean value $\Theta_{d}=48.7\mu as$ (blue dashed lines on top panels of Fig. 7 of the angular diameter of the supermassive black hole Sgr A*. It is clear from Fig. 7 that the values of the BH parameters $r_{s}$ and $\rho_{s}$ should be relatively high to coincide with the EHT observational data for Sgr A *.
@@
In addition, we performed the same analysis for the measured value of the angular diameter of SMBH M87* (Event Horizon Telescope Collaboration et al. [2019]) in the bottom panel of Fig. 7. Again, we present the lower bound $\Theta_{d}=39\mu as$ (solid black lines in the bottom graphics of Fig. 7), and mean value $\Theta_{d}=42$ (blue dashed lines in the bottom panels of Fig. 7) of the angular diameter of the supermassive black hole M87*. Similarly, the characteristic scale $r_{s}$ and the characteristic density $\rho_{s}$ of the rotating BH in the dark matter halo have to be relatively high to coincide with the EHT observational data for M87*.
5.1 Parameter estimation for a rotating black hole in Dehnen-type DM halo through MCMC analysis
In this section we constrain DM parameter $\rho_{s}$, $r_{s}$ and BH mass $M$, BH spin $a$ employing MCMC analysis and using observed data for $SgrA^{*}$ from EHT collaboration (Event Horizon Telescope Collaboration et al. [2019]), observed data for $SgrA^{*}$ from Gravity collaboration (Abuter et al. [2022]), observed data for $M87^{*}$ from EHT collaboration (Akiyama and et al. (2022) [Event Horizon Telescope Collaboration]).
We start with posterior distribution:
$P\left(\Theta|\mathcal{D},\mathcal{M}\right)=\frac{\mathcal{L}\left(\Theta|\mathcal{D},\mathcal{M}\right)\pi\left(\Theta|\mathcal{M}\right)}{P\left(\mathcal{D},\mathcal{M}\right)}\,$, (40)
where the likelihood function $L\left(\Theta|\mathcal{D},\mathcal{M}\right)$ tells us how well the parameters $\Theta$ explain the data $\mathcal{D}$ according to model $\mathcal{M}$:
$\mathcal{L}=-\frac{1}{2}\Sigma_{i}\left(\frac{\theta_{obs.}^{i}-\theta_{th.}^{i}}{\sigma_{i}}\right)^{2}\,$, (41)
in which $\theta_{obs.}$ is the observed value of the angular shadow size of the BHs, $\theta_{th.}$ is the theoretical value (39) of the angular shadow size of the BHs.
Figure 8: MCMC for Sgr A* from EHT (top left), Sgr A* from Gravity (top right), M 87* from EHT (bottom).
Then we set the priors to be Gaussian priors, with boundary condition:
$\pi(\Theta_{i})\sim\exp{\left[\left(\frac{\Theta_{i}-\Theta_{0,i}}{\sigma_{i}}\right)^{2}\right]},$ (42)
in which $\sigma_{i}$ is the standard deviation. Subsequently, numerical values for the priors of the parameters of rotating BH in DM halo used in our analysis is given in Table (1).
$SgrA^{*}$ from EHT $SgrA^{*}$ from Gravity $M87^{*}$ from EHT
$\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$
[…]
$r_{s}[10^{17}M_{\odot}]$ 0.8 0.15 0.8 0.2 1.6 0.21
$\rho_{s}[1/10^{19}M_{\odot}^{2}]$ 0.255 0.005 0.0326 0.005 0.004 0.0007
$a[M]$ 0.9 0.15 0.9 0.16 1 0.1
Table 1: Gaussian prior on rotating BHs in Dehnen-type DM halo from angular shadow size constraints.
Following, we constrain DM halo parameters $\rho_{s}$, $r_{s}$ and BH mass $M$, BH spin $a$ employing MCMC analysis and show visually in Fig.(8). Also, to be more informative we give numerical values of best-fit values for the parameters in Table(2).
$SgrA^{*}$ from EHT $SgrA^{*}$ from Gravity $M87^{*}$ from EHT
$M(\times 10^{6}M_{\odot})$ $4.057^{+0.563}_{-0.528}$ $4.087^{+0.605}_{-0.537}$ $6.431^{+0.438}_{-0.442}\times 10^{3}$
$r_{s}[10^{16}M_{\odot}]$ $0.784^{+0.145}_{-0.121}$ $0.802^{+0.178}_{-0.132}$ $1.6^{+0.210}_{-0.208}$
$\rho_{s}[1/10^{19}M_{\odot}^{2}]$ $025^{+0.005}_{-0.005} 0.0$$33^{+0.005}_{-0.005} 0.0$$004^{+0.001}_{-0.001}$
$a[M]$ $0.899^{+0.147}_{-0.146}$ $0.9^{+0.149}_{-0.150}$ $1^{+0.100}_{-0.099}$
Table 2: Best-fit rotating BHs in DM halo parameters derived from angular shadow size.
Our derived best-fit parameters for the dark matter halo show well agreement with values reported in prior studies (see for example, Cautun et al. [2020]) as:
$\rho_{s}\sim\frac{10^{-2}}{(10^{19}M_{\odot})^{2}}\frac{c^{6}}{G^{3}}\sim 10^{-24}g/cm^{3}\,$, (43)
$r_{s}\sim 10^{17}M_{\odot}\frac{G}{c^{2}}\sim 10^{22}cm\sim 10\text{kpc}\,.$
6 Thin accretion disk
Now we aim to get a simulation of a geometrically thin infinite accretion disk in a rotating BH surrounded by a DM halo spacetime (8). To achieve our intention, we have modified the open source Gyoto ray-tracing code (Vincent et al. [2011]) to adapt the code to our obtained spacetime (8).
Finding innermost stable circular orbits (ISCO) is an important step to analyze the thin accretion disk of the rotating BH in DM halo (Marck [1996]). To do it we start with expressing the effective potential of the massive particle in the vicinity of a rotating BH surrounded by Dehnen-type DM halo (Khan et al. [2024]):
$V_{m.eff.}=\frac{g_{t\phi}}{g_{\phi\phi}}l+\sqrt{\left(\frac{g_{t\phi}^{2}}{g_{\phi\phi}}-g_{tt}\right)\left(1+g^{\phi\phi}l^{2}\right)},$ (44)
where $l=\frac{L}{m}$ is the specific angular momentum of the test particle. Then the ISCO position can be found using the condition:
$V_{m.eff.}=\frac{\partial V_{m.eff}}{\partial r}=\frac{\partial^{2}V_{m.eff}}{\partial r^{2}}=0.$ (45)
Figure 9: The dependence of the ISCO radius on the parameters of the DM halo $\rho_{s}$ and $r_{s}$ for different cases.
Then, solving Eq.(45) numerically, we have plotted how the ISCO radius depends on the DM halo parameters $\rho_{s}$ and $r_{s}$ for different cases in Fig.9. It is obvious from Fig.9 that increasing both DM halo parameters $\rho_{s}$ and $r_{s}$ leads to the enlargement of the value of the ISCO radius $r_{ISCO}$ since the DM halo causes the mass of the BH-DM system.
The flux of electromagnetic radiation can be expressed as (Uktamov et al. [2025b], Alloqulov et al. [2024], Alloqulov and Shaymatov [2024], Alloqulov et al. [2025a, b]):
$\mathcal{F}(r)=-\frac{\dot{M_{0}}}{4\pi}\frac{\Omega_{,r}}{\sqrt{-g}\left(E-\Omega L\right)^{2}}\int_{r_{ISCO}}^{r}\left(E-\Omega L\right)L_{,r}dr,$ (46)
here $\dot{M_{0}}$ is the accretion rate, $g$ is the determinant of the three-dimensional subspace.
Then, in Fig.10 we have plotted the image thin accretion disk in a rotating BH surrounded by a Dehnen-type DM halo for different values of the DM parameters $\rho_{s}$ and $r_{s}$ for the angle of inclination $\theta_{0}=\frac{\pi}{10}$. One can notice by comparing images of the thin accretion disks of the obtained rotating BHs with thin accretion disks of Kerr and Schwarzschild BHs in Fig.10, that increasing both DM halo parameters $\rho_{s}$ and $r_{s}$ causes decreasing the value of the radiation flux as these parameters $\rho_{s}$, $r_{s}$ lead to growing the mass of the BH-DM system, so extending gravity around rotating BH in the DM halo.
Figure 10: The images of the thin accretion disk in Schwarzschild BH, Kerr BH and rotating BH immersed in Dehnen-type DM halo spacetime with the observational angle $\theta_{0}=\frac{\pi}{10}$.
7 Conclusion
@@
The Event Horizon Telescope (EHT) observations of the shadows of supermassive black holes, like Sgr A* at the Milky Way’s center and M87* in the nearby Messier 87 galaxy, provide an exceptional and natural laboratory for studying the gravitational properties of spacetime geometry around black holes, allowing tests in both strong and weak gravitational field regimes. In this work, we have generated a rotating BH space-time in a dark matter halo by applying the NJA method to the static spacetime metric and the components of the obtained metric tensor are provided in (8). Then the dependence of the horizon radius on the BH spin $a$ is shown in Fig. 2 for different values of the characteristic density $\rho_{s}$ and the characteristic scale factor $r_{s}$. It is shown that increasing spacetime parameters $r_{s}$ and $\rho_{s}$ causes an enlargement of the horizon radius of the rotating BH in a dark matter halo.
@@
We have also studied the shadow cast by the spinning axially symmetric black hole surrounded by a Dehnen-type dark matter halo. Using the Hamilton-Jacobi equation, we found photon motion in the vicinity of the rotating BH in a dark matter halo, which enabled us to find the photon sphere $r_{ph}$ and the shadow radius. The results obtained showed that the presence of the dark-matter halo causes an enlarged shadow as presented in Fig. 3.
@@
Additionally, we have analyzed the photon sphere and shadow radius of the rotating BH in a dark matter halo by giving numerical values of the photon sphere and shadow radius in Fig. 4.
@@
To strengthen our pure theoretical analysis, we got constraints for BH parameters $\rho_{s}$ and $r_{s}$ from EHT observational data for spacetime parameters of supermassive black holes SgrA* and M87* in Sec. 5. Our provided analysis for the angular diameter $\Theta_{d}$ demonstrates that the characteristic density $\rho_{s}$ and the characteristic scale factor $r_{s}$ of the rotating BH in the Dehnen-type dark matter halo should be relatively high to fit the EHT observed data of the SMBHs images (look to Fig.7).
One may conclude from the performed study that the influence of a Dehnen-type dark matter halo in the environment of supermassive black holes Sgr A* and M87* can not be excluded. To be more precise, we perform MCMC analysis to constrain DM parameters $\rho_{s}$, $r_{s}$ and BH mass $M$, BH spin $a$ and it is presented that obtained best-fit values for the DM parameters $\rho_{s}$ and $r_{s}$ are well agreement with values showed in prior studies.
Finally, in Sec. 6 we have analyzed thin accretion disk in a new BH surrounded by a DM halo and give an image of the thin accretion disk in Fig.10. It was found that increasing DM parameters $\rho_{s}$ and $r_{s}$ causes a growing mass of the BH-DM system so leading to a decrease in the value of the electromagnetic radiation.
8 Acknowledgement